electric machinery

chapter dc motor

qus=

200V DC series motor has Ra=0.5Ω and Rf=0.2Ω. the current input is 20 A, and the speed is 1000 r/min. if a resistor 0.2 Ω is joined in parallel with the filed winding, find the new speed of the motor for an input current of 20 A. Assume linear magnetization curve
Answer #1


case i) Ra=0.52 Rf=o.ar mm + 26 Eb २००। Vt=200, I a=204 N = 1000rpm Ebi 200 – 20 (0.2 +0.5) 200 -14 = 186 volts Emp equation cas case (2) Re- 0.2 Ra=0.5 Im + 28 Rp = 0.2 T Eba SOOV Eba= 200 - (0.2|10-2 +0.5) 20 - 200 - colto.520 = 200-12 . 188 voltsin case ü flux = 0 if 20 A in case (2) 10A flux - If ola (- : linear magnetisation) given back emf equation of DC motor Ebe k